# Critique-Coder: Enhancing Coder Models by Critique Reinforcement Learning

### v1

@@ 2.2 Dataset Construction
Critique Data Generation.  Figure 3 illustrates the critique data generation workflow. For each problem, we prompt Qwen3-Coder-30B-A3B-Instruct (Yang et al., 2025) to generate outputs, from which we extract code blocks as candidate solutions. Empty code blocks are discarded to ensure that only valid programs are retained for evaluation. Each candidate solution is then executed on the test cases from the filtered dataset, and its pass rate is computed to determine its judgment. A practical challenge is that certain test cases exhibit excessively long execution times, which may cause timeouts and lead to the erroneous classification of correct solutions as failures. To relax this constraint, we adopt a pass rate threshold of 80%: candidate solutions are labeled as True if their test pass rate exceeds this percentage, and as False otherwise.
Hybrid Data Integration. In our case, training exclusively on critique-oriented data biases the model toward overly focusing on evaluative feedback rather than task-oriented solutions, thereby hindering its ability to directly generate answers in evaluation tasks. To mitigate this issue, we construct a hybrid dataset that combines both CRL and standard RL data. This approach balances the characteristics of both data types and exposes the model to a diverse range of scenarios during training. Concretely, we randomly assign 20% of the data from the dataset to be CRL data, with the remaining 80% consisting of standard RL data. Such a configuration not only mitigates the risk of format shift but also improves the robustness of the learning process by jointly exploiting the complementary advantages of both CRL and RL training paradigms.
@@ 3.2 Evaluation Setup
For the sampling configuration, we follow the thinking mode settings reported in the original Qwen3 paper (Yang et al., 2025), with a temperature of 0.6, a top-p of 0.95, a top-k of 20, and a maximum output length of 32,768 tokens. The same configuration is applied consistently across all evaluation tasks. For LiveCodeBench specifically, we adopt the official evaluation prompt used in Qwen3 thinking mode. We also compare with some existing strong coding models like DeepSeek-R1-distill-14B (Guo et al., 2025), DeepCoder (Luo et al., 2025), DeepSeek-V2.5 (DeepSeek-AI, 2024), and GPT-o1 (Jaech et al., 2024b) with high reasoning effort.
@@ 3.3 Main Results
Compared with the base models, Critique-Coder leads to consistent and notable improvements across benchmarks of varying difficulty levels. On Qwen3-4B, for example, the LiveCodeBench score of Critique-Coder rises from 54.2 to 59.0, a gain of +4.8, surpassing the larger Qwen3-8B baseline by +1.5 points. On the Aider-Polyglot benchmark, which consists of multiple programming languages, Critique-Coder still demonstrates strong performance, achieving a +7.2 improvement from Qwen3-8B baseline, despite being trained solely on Python data using CRL. These results indicate the effectiveness of our algorithm.
@@
Our 4B and 8B models are also highly competent in their absolute performance. Critique-Coder-4B can beat the DeepCoder-14B (Luo et al., 2025) significantly across the board despite having only 28% of its parameters. Critique-Coder-8B can beat other strong models and only lags behind GPT-o1 (Jaech et al., 2024b) on Aide-Polyglot, which is mainly due to not optimizing other languages beyond Python. On EvalPlus, Critique-Coder-4B scores an impressive 86.5, just behind GPT-o1, while Critique-Coder-8B takes it a step further with an even higher score of 87.7.
@@ 3.4 Transferable Reasoning Ability
To examine whether the critique and reasoning abilities learned by Critique-Coder extend beyond coding tasks, we further evaluate the model on the BIG-Bench Extra Hard (BBEH) logic reasoning benchmark (Kazemi et al., 2025). As shown in Table 4, Critique-Coder achieves consistent improvements over both the baseline Qwen3-4B and its RL-trained variant across all four reasoning subtasks. In particular, it outperforms the Qwen3-4B-RL model by +4.5 points on BoardgameQA, leading to an overall average gain of +4.0 points, and surpasses the base model by +6.1 points on average. These results indicate that the critique-enhanced training paradigm not only improves performance in coding benchmarks but also strengthens general reasoning capabilities.
@@ 3.5 Test Time Scaling
We further evaluate Critique-Coder-4B with sequence test-time scaling (budget forcing) (Muennighoff et al., 2025), which extends the reasoning length at inference. On LiveCodeBench (v5), removing the constraint on reasoning tokens enables the model to achieve a score of 62.0 after four iterations, representing a +3.0 improvement compared to the setting without test-time scaling, as reported in Figure 5. This result is highly valuable to show that our 4B model can even beat other much larger baselines like DeepCoder-14B.
@@ 3.6 Ablation Study
To examine how the proportion of CRL data affects training, we perform an ablation study by varying the percentage of RL data replaced with CRL data, while keeping the overall dataset size fixed. As shown in Table 5, introducing a moderate amount of CRL data consistently improves performance over the pure RL setting. In particular, using 20% CRL data yields the best overall results, surpassing both the baseline (0%) and higher proportions (50% and 100%) across most benchmarks. Notably, while 50% CRL still maintains comparable performance, fully replacing RL data with CRL leads to clear degradation, suggesting that CRL is most effective as a complement to RL rather than a complete substitute. These findings highlight the importance of balancing RL and CRL data, with a small proportion of CRL providing the most robust gains. Otherwise, there will be a mismatch between training outputs and inference behavior during evaluation, as training solely with CRL drives the model to focus excessively on judgment while neglecting solution generation.
@@ 3.7 Limitations in Self-Critique
Although incorporating CRL enhances the model’s reasoning and critique abilities, it does not exhibit self-critique capability. To explore this limitation, we implemented critique-based parallel test-time scaling on Critique-Coder. Specifically, on LiveCodeBench v5, the Critique-Coder-4B was prompted to generate 10 candidate solutions for each question. These solutions, along with their corresponding question, were subsequently fed back into the model for critique, with each solution being critiqued for 64 samples. We then attempted to identify the best solution by leveraging these critiques. The solutions were scored based on the number of True critiques they received. If multiple solutions received identical scores, the solution with the shortest critique thinking token was selected. However, this approach did not yield performance improvements, indicating that the model lacks genuine self-critical ability.

### v2

@@ 2 Method
This section introduces the training process of Critique-Coder, which combines standard reinforcement learning and critique reinforcement learning for code generation. We first present the problem formulation and optimization method, followed by dataset construction and the overall training procedure.
@@ 2.2 Dataset Construction
Critique Data Generation.  Figure 3 illustrates the critique data generation workflow. For each problem, we prompt Qwen3-Coder-30B-A3B-Instruct (Yang et al., 2025) to generate outputs, from which we extract code blocks as candidate solutions. Despite its larger parameter scale, Qwen3-Coder-30B-A3B-Instruct lacks explicit reasoning capabilities, and its effective problem-solving performance is not stronger than that of smaller reasoning-enabled models such as Qwen3-4B (Thinking). Empirically, Qwen3-Coder-30B-A3B-Instruct achieves a pass@1 of 32.20% on our dataset, compared to 43.72% for Qwen3-4B (Thinking). Owing to this relatively weak performance, the critique data generation process does not constitute a distillation setup, as the data are not produced by a stronger teacher model. Empty code blocks are discarded to ensure that only valid programs are retained for evaluation. Each candidate solution is then executed on the test cases from the filtered dataset, and its pass rate is computed to determine its judgment. A practical challenge is that certain test cases exhibit excessively long execution times, which may cause timeouts and lead to the erroneous classification of correct solutions as failures. To relax this constraint, we adopt a pass rate threshold of 80%: candidate solutions are labeled as True if their test pass rate exceeds this percentage, and as False otherwise.
Hybrid Data Integration. In our case, training exclusively on critique-oriented data biases the model toward evaluative behaviors, encouraging it to focus on judging or analyzing candidate solutions rather than directly producing task-oriented outputs, which in turn degrades its ability to generate complete answers in evaluation tasks. Standard RL data, by contrast, explicitly reinforces end-to-end solution generation through task-level rewards but does not expose the model to reflective or critique-based reasoning patterns. To mitigate this imbalance, we construct a hybrid dataset that combines both CRL and standard RL data, allowing the model to jointly learn evaluative judgment and direct solution synthesis within a unified training process. Concretely, we randomly assign 20% of the data from the dataset to be CRL data, with the remaining 80% consisting of standard RL data. Such a configuration not only mitigates the risk of format shift caused by overexposure to critique-style supervision but also improves the robustness of the learning process by exploiting the complementary strengths of CRL and RL training paradigms.
@@ 3 Experiment
This section presents a comprehensive empirical evaluation to demonstrate the effectiveness of Critique-Coder. We report training and evaluation settings, compare against baseline and RL-trained models on multiple coding benchmarks, and further analyze transferability, test-time scaling, and ablation results.
@@ 3.2 Evaluation Setup
For the sampling configuration, we follow the thinking mode settings reported in the original Qwen3 paper (Yang et al., 2025), with a temperature of 0.6, a top-p of 0.95, a top-k of 20, and a maximum output length of 32,768 tokens. The same configuration is applied consistently across all evaluation tasks. For LiveCodeBench, we adopt the official evaluation prompt used in Qwen3 thinking mode to ensure consistency. We further compare our model with several strong coding baselines, including DeepSeek-R1-distill-14B (Guo et al., 2025), DeepCoder (Luo et al., 2025), DeepSeek-V2.5 (DeepSeek-AI, 2024), and GPT-o1 (Jaech et al., 2024b) with high reasoning effort, and report results from their official publications or recommended evaluation settings.
@@ 3.3 Main Results
Compared with the base models, Critique-Coder leads to consistent and notable improvements across benchmarks of varying difficulty levels and evaluation protocols. On Qwen3-4B, for example, the LiveCodeBench score of Critique-Coder increases from 54.2 to 59.0, yielding a substantial gain of +4.8 and even surpassing the larger Qwen3-8B baseline by +1.5 points. On the Aider-Polyglot benchmark, which evaluates multi-language programming ability, Critique-Coder achieves a +7.2 improvement over the Qwen3-8B baseline, despite being trained solely on Python data using CRL. Importantly, these gains are consistently reflected across all evaluated benchmarks and model scales, and also translate into higher overall average scores (+2.4 for 4B and +3.5 for 8B), indicating that the improvements are broad-spectrum rather than limited to specific tasks or datasets.
@@
Despite their relatively small parameter sizes, our models demonstrate strong absolute performance when compared with frontier systems. Critique-Coder-4B significantly outperforms DeepCoder-14B (Luo et al., 2025) across all evaluated benchmarks while using only 28% of its parameters, highlighting the parameter efficiency of our approach. Critique-Coder-8B remains competitive with other strong models and only lags behind GPT-o1 (Jaech et al., 2024b) on Aider-Polyglot, which can be largely attributed to its lack of explicit optimization for non-Python languages. On EvalPlus, Critique-Coder-4B achieves an impressive score of 86.5, closely trailing GPT-o1, while Critique-Coder-8B further improves this result to 87.7, demonstrating that our method scales effectively with model size.
@@ 3.4 Transferable Reasoning Ability
To examine whether the critique and reasoning abilities learned by Critique-Coder extend beyond coding tasks, we evaluate the model on the BIG-Bench Extra Hard (BBEH) logic reasoning benchmark (Kazemi et al., 2025). As shown in Table 4, Critique-Coder consistently outperforms both the baseline Qwen3-4B and its RL-trained variant across all four reasoning subtasks, achieving an average improvement of +6.1 points over the base model. The gains are especially pronounced on more structured and constraint-intensive tasks such as Time Arithmetic (+8.0) and Zebra Puzzles (+7.5), which require multi-step logical consistency rather than surface-level pattern matching. Moreover, despite being optimized with reinforcement learning, Qwen3-4B-RL shows only limited improvements, whereas Critique-Coder achieves an additional +4.0 average gain and outperforms it by +4.5 points on BoardgameQA. Notably, we use the same inference configuration for all models, suggesting that the gains are driven by critique-enhanced training rather than evaluation-time settings. Overall, these results indicate that critique-enhanced training provides additional benefits over standard RL optimization and effectively improves reasoning abilities that transfer beyond coding tasks.
@@ 3.5 Test Time Scaling
We further evaluate Critique-Coder-4B with sequence test-time scaling (budget forcing) (Muennighoff et al., 2025), which relaxes the constraint on reasoning tokens at inference to enable longer reasoning trajectories. As shown in Figure 5, increasing the test-time computation budget leads to a consistent and monotonic improvement on LiveCodeBench (v5): starting from the default setting without test-time scaling (approximately 13k tokens per question and 59.0 accuracy), the model steadily improves as the number of scaling iterations increases, reaching an accuracy of 62.0 after four iterations and achieving a +3.0 absolute gain over the no-scaling baseline. Further increases in the reasoning budget yield marginal but stable improvements, indicating diminishing returns. Overall, these results demonstrate that test-time scaling is an effective mechanism for enhancing the performance of a compact 4B code model without modifying its parameters. We note that this improvement is specific to sequence-based test-time scaling; in contrast, parallel test-time scaling does not yield gains, which we analyze in detail in subsection 3.7.
@@ 3.6 Ablation Study
To study the impact of the CRL data ratio, we conduct an ablation study by replacing different proportions of RL data with CRL data while keeping the total amount of training data and optimization settings fixed. All training samples are derived from the same filtered rStar-Coder seed dataset, and CRL data are obtained by converting a subset of the original RL samples following the same procedure. As shown in Table 5, introducing a moderate amount of CRL data consistently improves performance over the pure RL baseline. Among all configurations, the model trained with 20% CRL data achieves the best overall results, increasing the average score from 45.7 to 47.2 (+1.5). This setting yields the strongest or tied-best performance on most benchmarks, including EvalPlus, BigCodeBench-I (Full), Aider-Polyglot, and LiveCodeBench v5, demonstrating that CRL can effectively complement RL training.
As the CRL ratio increases further, the performance gains gradually diminish. Although the 50% CRL setting remains competitive on average, it shows degradation on more challenging subsets such as BigCodeBench-I (Hard). When RL data are completely replaced by CRL data, performance drops significantly across multiple benchmarks, resulting in an average score (44.5) lower than the baseline. These results suggest that CRL should be used as a complementary signal rather than a replacement for RL data. We hypothesize that excessive reliance on CRL may bias the model toward judgment-oriented behaviors, introducing a mismatch with inference-time requirements that emphasize long-horizon solution generation, particularly on harder benchmarks. Overall, maintaining an appropriate balance between RL and CRL data is crucial, with 20% CRL providing the most favorable trade-off in our experiments.
@@ 3.7 Limitations in Self-Critique
Although incorporating CRL substantially improves the model’s reasoning and critique-related capabilities during generation, it does not enable reliable post-hoc self-critique when only the final answer is available. To examine this limitation, we implemented critique-based parallel test-time scaling on Critique-Coder. Specifically, on LiveCodeBench v5, Critique-Coder-4B was prompted to generate 10 candidate solutions per problem. Each solution, together with the original question, was then fed back into the model for critique, with 64 critique samples generated per solution. Candidate solutions were scored by the number of True critiques they received, with ties broken by selecting the solution with the shortest critique thinking token length. This procedure did not lead to performance improvements.
This outcome is not specific to our method and is consistent with prior findings (Huang et al., 2023; Valmeekam et al., 2023) showing that large language models cannot reliably evaluate or refine their own final outputs without access to intermediate reasoning trajectories or additional external signals. Importantly, this result does not contradict the effectiveness of CRL. The performance gains from CRL do not arise from post-hoc self-critique on completed answers; instead, CRL improves performance by internalizing critique-based judgment during training, enabling more thorough reasoning and refinement before the final answer is produced. Consequently, once the reasoning process is removed and only the final output is provided, the model naturally fails to further critique or improve it, in line with established limitations in the literature.